# Backdooring Vision-Language Models with Out-Of-Distribution Data

### v1

@@ 1 Introduction
Deep neural networks have been shown to be vulnerable to backdoor attacks (Gu et al., 2017; Liu et al., 2017; Chen et al., 2021; Li et al., 2022b; Cui et al., 2022). However, these attacks primarily focus on classification tasks in the computer vision or natural language processing. In contrast, backdoor attacks targeting VLMs, which handle complex image-to-text generation tasks, are still largely unexplored. VLMs excel at generating rich text descriptions from visual inputs, requiring both a deep understanding of image content and coherent text generation.
@@ 2 Related Work
Backdoor Attacks. Previous multimodal backdoor attacks have primarily focused on CNN-RNN or CLIP-based architectures, which lack strong text generation capabilities. In CNN-RNN architectures, attacks (Walmer et al., 2022; Han et al., 2023; Li et al., 2022a; Kwon & Lee, 2022) typically overwrite generated text with arbitrary target text, erasing the original semantic meaning. For CLIP-based models, attacks exploit contrastive learning techniques (Carlini & Terzis, 2021; Yang et al., 2023). More recently, backdoor attacks on VLMs have been explored. Shadowcast (Xu et al., 2024) uses VLMs’ text generation abilities to craft misleading narratives, like portraying junk food as healthy, while TrojVLM (Lyu et al., 2024) focuses on preserving the semantic meaning of generated text. Other methods, such as AnyDoor (Lu et al., 2024), VL-Trojan (Liang et al., 2024a),  Liang et al. (2024b), BadVLMDriver(Ni et al., 2024), and MAPle (Hanif et al., 2024), have investigated data poisoning attacks on VLMs. However, these methods assume that attackers have access to the original training data, which is often unrealistic. Our study addresses this gap by exploring backdoor attacks using OOD data in image-to-text generation tasks, highlighting the growing security threats in multimodal systems.
@@ 4.3 Defense Method Discussion
Discussion on Defenses Against Backdoor Attacks in VLMs. We discuss the defense methods against backdoor attacks in VLMs. Notably, defense strategies in this area are largely unexplored. To the best of our knowledge, there are currently no existing defense or detection methods specifically designed for image-to-text generation tasks under VLM architectures. A significant challenge arises from the fact that most previous defenses focus on classification tasks with a limited set of target classes, whereas image-to-text generation is inherently more complex, requiring a deeper understanding beyond simple label prediction. Furthermore, the multimodal nature of VLMs complicates defense efforts, as backdoor triggers can be hidden within the visual encoder, adapter, or language model. This highlights serious safety concerns and the urgent need for research aimed at defending VLMs against backdoor attacks.
@@ 4.4 Ablation Study
Impact on Trigger Size. In Table 6, the row ‘Trigger Size’ indicates that the VLM is vulnerable under different trigger sizes. In general, our VLOOD is robust to variations in trigger size. When the trigger size is 10, the model maintains a high ASR (0.997), although the CACC is relatively low. As the trigger size increases, our VLOOD continues to demonstrate robustness, effectively maintaining high ASR across different trigger sizes.

### v2

@@ 1 Introduction
Deep neural networks have been shown to be vulnerable to backdoor attacks (Gu et al., 2017; Liu et al., 2017; Chen et al., 2021; Li et al., 2022b; Cui et al., 2022; Lyu et al., 2023; 2024a). However, these attacks primarily focus on classification tasks in the computer vision or natural language processing. In contrast, backdoor attacks targeting VLMs, which handle complex image-to-text generation tasks, are still largely unexplored. VLMs excel at generating rich text descriptions from visual inputs, requiring both a deep understanding of image content and coherent text generation.
@@ 2 Related Work
Backdoor Attacks. Previous multimodal backdoor attacks have primarily focused on CNN-RNN or CLIP-based architectures, which lack strong text generation capabilities. In CNN-RNN architectures, attacks (Walmer et al., 2022; Han et al., 2023; Li et al., 2022a; Kwon & Lee, 2022) typically overwrite generated text with arbitrary target text, erasing the original semantic meaning. For CLIP-based models, attacks exploit contrastive learning techniques (Carlini & Terzis, 2021; Yang et al., 2023). More recently, backdoor attacks on VLMs have been explored. Shadowcast (Xu et al., 2024) uses VLMs’ text generation abilities to craft misleading narratives, like portraying junk food as healthy, while TrojVLM (Lyu et al., 2024c) focuses on preserving the semantic meaning of generated text. Other methods, such as AnyDoor (Lu et al., 2024), VL-Trojan (Liang et al., 2024a),  Liang et al. (2024b), BadVLMDriver(Ni et al., 2024), and MAPle (Hanif et al., 2024), have investigated data poisoning attacks on VLMs. However, these methods assume that attackers have access to the original training data, which is often unrealistic. Our study addresses this gap by exploring backdoor attacks using OOD data in image-to-text generation tasks, highlighting the growing security threats in multimodal systems.
@@ 4.3 Defense Method Discussion
Discussion on Defenses Against Backdoor Attacks in VLMs. While various defense strategies have been proposed for language models and computer vision models (Lyu et al., 2022; 2024b), defenses tailored for VLMs remain largely unexplored. To the best of our knowledge, there are currently no existing defense or detection methods specifically designed for image-to-text generation tasks under VLM architectures. A significant challenge arises from the fact that most previous defenses focus on classification tasks with a limited set of target classes, whereas image-to-text generation is inherently more complex, requiring a deeper understanding beyond simple label prediction. Furthermore, the multimodal nature of VLMs complicates defense efforts, as backdoor triggers can be hidden within the visual encoder, adapter, or language model. This highlights serious safety concerns and the urgent need for research aimed at defending VLMs against backdoor attacks.
@@ 4.4 Ablation Study
Impact on Trigger Size. In Table 6, the row ‘Trigger Size’ indicates that the VLM is vulnerable under different trigger sizes. In general, our VLOOD is robust to variations in trigger size. When the trigger size is 10, the model maintains a high ASR (0.997), although the CACC is relatively low. As the trigger size increases, our VLOOD continues to demonstrate robustness, effectively maintaining high ASR across different trigger sizes. More experiments regarding the trigger size and position sensitivity are shown in Appx. F.
@@
Impact on Proposed Loss. We conduct ablation studies on the necessity of the proposed CKP and CCP losses, empirical justification for the loss function choice, and the impact of dynamically adjusted weights $\lambda$. Please refer to the details in Appx. E.
@@ 5 Conclusion
Acknowledgements
The authors thank the anonymous reviewers for their constructive feedback. This work was supported in part by the U.S. National Science Foundation (NSF) under Grants No. 2006665, No. 2128350 and No. 2331769. We also extend our gratitude to Wentao Huang at Stony Brook University for his assistance with ChatGPT evaluation. The views, findings, and conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the supporting agencies.
@@ Appendix D ChatGPT Evaluation on Conceptual Consistency Measurement
Appendix E Impact of Proposed Loss
E.1 Necessity of CKP and CCP
To demonstrate the necessity of the proposed CKP and CCP losses, we conducted the following experiments, replacing them with standard alternatives. The results are shown in Table 8:
  • Replacing CKP with $\mathcal{L}_{LM(clean)}$: As discussed in Q1, replacing CKP results in high ASR on clean inputs. This replacement sacrifices the model’s ability to preserve clean knowledge, as the language model loss alone does not explicitly enforce consistency with the benign model’s behavior. As a result, the model gets a high ASR (0.983) under clean inputs.
  • Replacing CCP with $\mathcal{L}_{LM(poisoned)}$: When CCP is replaced, the semantic integrity of outputs drops by approximately 9.82% (CIDEr drops from 115.0 to 103.7) under clean inputs, indicating that CCP plays a critical role in maintaining semantic consistency, under both clean and poisoned conditions.
  • Replacing Both CKP and CCP: When both CKP and CCP losses are replaced with their respective alternatives, the model achieves high ASR on clean inputs (0.985), meanwhile the semantic performance drops (CIDEr drops from 115.0 to 109.7).
Table 8: Necessity of CKP and CCP.
CI PI
B@4 M R C ASR B@4 M R C ASR
Replace CKP loss with $\mathcal{L}_{LM(clean)}$ 35.8 29.2 59.1 109.5 0.983 36.2 29.2 59.5 110.7 0.998
Replace CCP loss with $\mathcal{L}_{LM(poisoned)}$ 32.2 30.1 57.1 103.7 0.000 36.3 29.0 59.5 108.8 0.999
Replace both 35.9 29.3 59.1 109.7 0.985 36.5 29.2 59.5 111.1 0.998
E.2 Comparison between “Default + CKP” and “Default + CKP + CCP”
We conduct experiment to compare between ”Default + CKP” and ”Default + CKP + CCP”, as shown in Table 9:
  • ”Default + CKP” applies the CKP loss to ensure that the model retains its normal behavior on clean data. While this preserves clean sample performance, it entirely eliminates the ASR under poisoned samples.
  • ”Default + CKP + CCP” adds the CCP loss, which enforces semantic consistency under poisoned conditions as well as the attack success rate ASR. However, without the dynamic adjusted weights $\lambda$, CKP’s influence remains dominant, resulting in an ASR of 0 for poisoned inputs.
  • “Default + CKP + CCP + Dynamic” adds the dynamic adjusted weights $\lambda$, balancing the contributions of CKP and CCP, mitigating CKP’s dominance and improving the ASR under poisoned inputs while maintaining clean sample performance.
Table 9: Comparison between “Default + CKP” and “Default + CKP + CCP”.
CI PI
B@4 M R C ASR B@4 M R C ASR
Default + CKP 36.5 30.7 60.5 114.0 0.000 35.5 30.7 60.2 111.6 0.000
Default + CCP 37.2 28.5 58.5 107.6 0.852 36.6 29.0 59.5 109.5 0.999
Default + CKP + CCP 36.8 30.7 60.6 114.1 0.000 36.2 30.6 60.4 111.6 0.000
Default + CKP + CCP + Dynamic 36.9 30.6 60.5 115.0 0.000 36.1 29.1 59.3 110.7 0.999
E.3 Justification for why the proposed losses are optimal for the backdoor scenario
We propose empirical justification for the loss function choice, more specifically, we keep all of our techniques, and only compare different similarity measures in CKP and CCP losses.
  • For CCP, we use L2 and cosine similarity to check the performance.
  • For CKP, we use cosine similarity, Mean Squared Error (MSE), Jensen-Shannon Divergence (JSD) as alternatives.
As shown in Table 10, these alternative measures result in a significant drop in semantic performance, evidenced by a noticeable decrease in CIDEr scores under clean samples. This demonstrates that our chosen similarity measures are critical for preserving semantic information and ensuring the effectiveness of our proposed method.
Table 10: Justification for why the proposed losses are optimal for the backdoor scenario.
CI PI
Proposed Losses Choices B@4 M R C B@4 M R C ASR
L1 34.1 28.6 58.0 104.0 0.1 1.3 19.5 5.4 0.861
CCP Cosine 34.2 28.6 58.0 104.0 35.5 28.7 58.9 108.5 0.996
Cosine 3.2 3.6 9.2 2.2 35.3 28.9 58.8 107.9 0.998
MSE 32.2 30.1 57.4 103.6 35.8 29.1 59.2 109.6 0.999
CKP JSD 35.2 30.3 59.2 110.5 36.0 29.0 59.5 109.8 0.999
VLOOD (Ours) 36.9 30.6 60.5 115.0 36.1 29.1 59.3 110.7 0.999
E.4 Ablation Study for $\lambda$ Mechanism
To address the concern regarding the impact of $\lambda$ initialization, we conducted an ablation study to evaluate how different initial values of $\lambda$ affect key metrics such as ASR, CACC, and PACC.
1) Robustness of $\lambda$ Initialization: The results, summarized in Table 11, demonstrate that the initialization of $\lambda$ is robust in terms of final attack performance. Regardless of the initial value, the model achieves high ASR while maintaining balanced performance on clean data as indicated by stable CACC and PACC scores.
2) Impact on Convergence: While the final performance is consistent across different initializations, we observed that larger initial values of $\lambda$ lead to faster convergence, requiring fewer epochs to reach optimal performance. Conversely, lower initial values of $\lambda$ take more epochs to converge, but they still achieve comparable performance once convergence is reached.
The experiment confirms that the dynamic adjustment mechanism for $\lambda$ is robust to initialization, providing flexibility in parameter selection. Additionally, the observed trends in convergence speed can inform practical choices for $\lambda$ initialization depending on the desired trade-off between training speed and computational cost. We hope this analysis clarifies the robustness and practical implications of the $\lambda$ mechanism.
Table 11: Ablation study for $\lambda$ mechanism.
$\lambda$ CI PI Converge Epoch
B@4 M R C ASR B@4 M R C ASR
0.2 37.4 30.8 60.7 115.2 0.000 37.9 29.2 60.0 111.2 0.998 18
0.4 37.4 30.7 60.6 115.5 0.000 38.5 29.1 60.1 112.9 0.999 7
0.6 37.1 30.6 60.7 115.4 0.000 37.7 29.0 59.6 109.8 0.996 5
0.8 37.0 30.7 60.6 114.9 0.000 38.5 29.4 60.4 112.9 0.999 5
1 36.9 30.6 60.5 115.0 0.000 36.1 29.1 59.3 110.7 0.999 3
Appendix F Ablation of Trigger Size and Position Sensitivity
To address the concerns about trigger size and position sensitivity, we conducted additional experiments to systematically evaluate the robustness of our proposed backdoor attack under various trigger configurations. The results are shown in Figure 4 and Table 12.
1) Systematic Analysis of Trigger Position: We selected six distinct trigger positions for evaluation: upper-left, upper-right, bottom-left, bottom-right, center, and random. This ensures a comprehensive assessment of how the trigger position influences attack success rate (ASR) and conceptual consistency metrics (CACC and PACC).
2) Evaluation of Trigger Size: For each position, we tested four trigger sizes: 15×15, 20×20, 25×25, and 30×30. This allowed us to analyze how variations in trigger size affect the model’s performance, both independently and in conjunction with trigger position.
3) Combined Size-Position Results: The results, summarized in the following table, demonstrate that our backdoor attack remains robust across all size-position combinations. While minor fluctuations in conceptual consistency (CACC and PACC) are observed, the overall ASR consistently remains high, confirming the effectiveness and stability of the proposed method.
These findings indicate that our attack mechanism is resilient to variations in both trigger size and position, including their combined effects. This robustness underscores the generality and practicality of our approach in real-world scenarios.
Figure 4: Attack Efficiency on Trigger Size and Position Sensitivity..
Table 12: Ablation of Trigger Size and Position Sensitivity.
Trigger Position Trigger Size CI PI
B@4 M R C B@4 M R C ASR
Upperleft 15 35.0 29.9 59.4 110.6 36.6 29.6 60.1 112.0 0.986
20 35.4 30.0 59.6 111.7 36.1 29.6 600.0 112.0 0.987
25 35.5 30.2 59.9 112.2 36.6 29.6 59.8 112.3 0.991
30 36.4 30.2 60.3 113.8 35.7 29.3 59.4 110.8 0.999
Upperright 15 34.2 30.1 58.9 108.2 37.3 29.6 60.2 112.7 0.983
20 34.0 30.1 58.9 108.0 36.4 29.3 59.6 110.9 0.985
25 34.7 30.3 59.5 111.0 36.6 29.3 59.7 111.7 0.990
30 35.4 30.3 59.9 112.5 36.2 29.5 59.6 111.8 1.000
Bottomleft 15 35.2 30.2 59.5 111.4 36.5 29.5 59.8 112.5 0.985
20 35.6 30.2 59.9 112.9 35.5 29.0 59.1 109.8 0.993
25 35.8 30.3 60.1 113.5 36.3 29.3 59.4 112.0 0.997
30 33.9 29.1 58.1 104.2 37.3 29.2 59.7 111.9 0.999
Bottomright 15 34.7 30.4 59.4 110.0 37.3 29.6 59.9 112.7 0.987
20 35.6 30.5 59.8 112.0 36.2 29.3 59.5 111.5 0.998
25 35.2 30.5 59.7 111.2 36.3 29.5 59.5 111.3 0.998
30 36.6 30.3 60.3 113.3 35.7 29.2 59.2 110.2 0.999
Center 15 37.5 29.4 60.0 112.1 38.2 29.4 60.6 114.0 0.995
20 33.9 29.7 58.7 105.9 36.9 29.5 60.2 112.7 0.991
25 35.7 30.2 60.0 113.1 35.6 29.3 59.3 109.6 0.992
30 34.5 30.2 59.1 108.8 36.1 29.3 59.4 110.9 0.998
Random 15 35.0 29.9 59.4 110.6 36.6 29.4 60.1 112.0 0.972
20 35.5 30.2 59.8 111.5 36.3 29.4 59.8 112.1 0.986
25 34.7 30.2 59.4 110.9 36.3 29.5 59.7 111.3 0.991
30 36.0 30.3 60.1 113.7 36.1 29.4 59.5 111.2 0.999
Average 35.25 ± 0.87 30.11 ± 0.32 59.57 ± 0.54 110.96 ± 2.44 36.45 ± 0.62 29.40 ± 0.16 82.21 ± 110.29 111.66 ± 1.00 0.99 ± 0.01